# Searching Latent Program Spaces

### v2

@@ 6 Conclusion
7 Acknowledgments and Disclosure of Funding
We thank Google’s TPU Research Cloud (TRC) for supporting this research. We are also grateful to Nathan Grinsztajn, Natasha Butt, Levi Lelis, and Jessica Hu for their feedback on the early versions of the paper. Matthew Macfarlane is supported by the LIFT-project 019.011, which is partly financed by the Dutch Research Council (NWO). We also thank reviewers for their detailed feedback which we have integrated into the final version of the paper.
@@ A.2 ARC-AGI
We evaluate LPN on the ARC-AGI benchmark by training it on the re-arc dataset [Hodel, 2024],a dataset designed to be in distribution relative to the ARC training set, previously used for training large language models and ARC specific transformers to solve ARC-AGI Li et al. [2024b], Akyürek et al. [2024]. Most previous works however expand beyond this dataset, usually synthetically adding additional programs [Butt et al., 2024], or crafted datasets. We restrict training to tasks from only this dataset with the goal of assessing generalization instead of closing a performance gap with more data. This also ensures LPN solely learns from core knowledge priors without data leakage from the evaluation set. The evaluation set is known to be significantly out of distribution and therefore represents a challenging generalization experiment. We train a 178M-parameter LPN with a 256-dim latent space for 100k steps (batch size 256) for 2 days on a TPU v4-32, see Appendix E for further details, in terms of data, this amounts to 51M I/O pairs. Leveraging insights from smaller scale experiments we make use of gradient in LPN during training. We train LPN in Grad 0 mode for 95k steps and then fine tune in Grad 1 for a further 5k steps. We outline our hyperparameter procedure for both LPN and TTT for ARC-AGI in Appendix E.
@@ B.1 LPN Training Extended
We provide an expanded version of Table 1 with additional ablations for both the training and inference axes.
@@
Table 4: Ablation study of LPN training and inference methods on the Pattern task. Rows represent training methods: LPN Grad [N] for N gradient steps, LPN Grad [N] ** with gradient flow through latent optimization (analogous to meta-learning), and Sample [N] for N random samples. Columns represent inference methods: Grad [N] for N gradient ascent steps and Sample [N] for random search with N samples. Training was performed for 20k steps with 3 seeds, with performance reported as mean (standard deviation) over 3 runs. Bold values indicate the best performance per inference method.
@@ B.2 Adapting Out-Of-Distribution
We provide expanded results for the out-of-distribution (OOD) experiments outlined in Section 5.4. We include results for varying levels of OOD starting with in-distribution, weak OOD, and strong OOD.
@@ B.3 Validating the Decoder
Training deep networks from scratch to solve ARC-like tasks has been challenging [Li et al., 2024c]. If it is the case that neural networks struggle even to learn to execute single programs this represents a significant bottleneck to training models from scratch on a broad distribution of programs. Therefore, before training LPN end-to-end, we conclusively show that our decoder architecture does not suffer from such a bottleneck, and can learn individual programs.
We show 5 of the 400 total tasks from the ARC-AGI training set, and for each of these tasks, we train a small LPN architecture of 800k parameters (except for the last task which required a bigger model with 8.7M parameters) on the corresponding task generator from re-arc [Hodel, 2024]. Specifically, we select the first five tasks from the arc-agi_training_challenges json file (007bbfb7, 00d62c1b, 017c7c7b, 025d127b, 045e512c) shown in figure Figure 5.
@@
We evaluate both the distribution of re-arc generated tasks on which it is trained and on the true task from the ARC-AGI training set in figure Figure 6. We show that for each task, the small LPN decoder-only model successfully learns individual programs and manages to solve the corresponding ARC-AGI task. Therefore, our model outperforms previously reported results in Li et al. [2024c], and concludes that our architecture does not suffer from a decoder bottleneck. Note that the encoder is not helpful in this experiment since the task is always the same. Our later results on ARC-AGI Section 5.6 take this a step further and show that we can learn a single transformer architecture capable of executing over 270 programs in the ARC training dataset.
@@ B.4 Analyzing the Latent Space
To validate that the encoder is learning programs in its latent space, we design an even simpler task with small 4x4 grids that have 2x2 patterns. We train a small LPN model until convergence with a latent space of dimension 2 to easily visualize it in figure Figure 7. Due to the simplicity of the task we train the model with mean training, i.e. no latent optimization. Because we are using a 2D Gaussian prior for the latent space, we can convert $\mathbb{R}^{2}$ to the unit square using the normal cumulative distribution function (CDF), and then plot on the unit square at coordinates $(x,y)$ the decoder’s output when conditioned by the latent $\textrm{CDF}(z)=(x,y)$, or equivalently, $z=(\textrm{PPF}(x),\textrm{PPF}(y))$ using the percent-point function (PPF). These results demonstrate the diversity and smoothness of the latent space, showing structure in terms of color patterns, which motivates performing gradient ascent latent optimization in more complex tasks. This shows that the latent space can encode a wide range of diversity in its latent space which is especially important for adapting to unseen patterns.
@@ B.5.1 Performance Across Training Sizes
Table 8: Performance (top-2 accuracy, percentage) for training size 11 on the pattern task.
@@ B.6 Measuring Floating Point Operations
The LPN Sampling results can be understood as ablation of LPN without gradient-based search, but only sampling multiple latents from the encoder computing the decoder likelihood for each, avoiding the computational cost of backpropagating through the decoder. However, this efficiency gain comes at a significant cost in performance, as shown in the pattern task results Table 1.
@@ B.7.1 Dataset
B.7.2 Results
Our experiments demonstrate that Test-Time Training (TTT) exhibits overfitting, as evidenced by their performance degradation when gradient fine tuning is performed. In contrast, LPN maintains robust performance without overfitting. Furthermore, incorporating a single gradient step during training enhances LPN performance marginally at higher inference gradients.
Inference (Avg $\pm$ Std)
Training Grad 0 Grad 10 Grad 100
In-Context 0.78 (0.01) - -
Test-Time Training 0.78 (0.01) 0.73 (0.01) 0.12 (0.01)
LPN Grad 0 0.76 (0.02) 0.83 (0.01) 0.81 (0.01)
LPN Grad 1 0.71 (0.00) 0.86 (0.01) 0.82 (0.02)
LPN Grad 2 0.63 (0.01) 0.85 (0.00) 0.80 (0.01)
LPN Grad 3 0.49 (0.04) 0.85 (0.00) 0.82 (0.02)
LPN Grad 1 ** 0.68 (0.01) 0.85 (0.00) 0.81 (0.01)
Table 9: Study of model performance across different inference gradients for methods: In-Context, Test-Time Training with learning rate of 1e-4, LPN Grad 0, LPN Grad 1, LPN Grad 2, LPN Grad 3 (all with learning rates of 1e-1), LPN Grad 1 **. Performance is averaged over 3 training runs, with standard deviation in parentheses.
@@ Appendix C LPN Algorithm
Below we outline two algorithms: first, LPN test-time inference (Algorithm 1) and its mechanism for performing inductive inference. Second, we provide the full algorithm for LPN during training (Algorithm 2).
@@
5:         Sample $z_{i}\sim q_{\phi}(z|x_{i},y_{i})$triangleright Using the reparameterization trick
@@
10:              $z_{i}^{\prime}\leftarrow z_{i}^{\prime}+\alpha\cdot\nabla_{z}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{n}\log p_{\theta}(y_{j}|x_{j},z)|_{z=z_{i}^{\prime}}$triangleright Optional stop-grad on the 2nd term
@@
14:     $\mathcal{L}\leftarrow\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}_{i}$triangleright Total loss for all pairs
@@ Appendix D Variational Inference
Latent Program Networks (LPNs), when operated in grad 0 mode, perform a form of variational inference. Updating the latent representation using gradients from the encoder constitutes semi-amortized variational inference [Kim et al., 2018, Marino et al., 2018].
In prior work, such as LEAPs Trivedi et al. [2021], it is assumed during training that the full representation of a program $f$ is observable. In this setting, variational inference on programs can be conducted using a standard Variational Autoencoder (VAE) restricted to the program space. LEAPs introduce a form of variational inference via an Execution Evidence Lower Bound (ELBO), where function reconstruction is based on correctly executing the function for a given input rather than reconstructing the full program representation. The Execution ELBO is defined as:
@@ Appendix E Hyperparameters
For the ARC-AGI results, we performed a hyperparameter search over the learning rate for test time adaptation for both test-time tuning and LPN. Since these methods operate in different spaces, they are unlikely to require similar parameters, making it fairer to tune this parameter independently for each baseline. We validated on a held-out test set of unseen problems from the RE-ARC dataset. We searched over learning rates ranging from $10^{-1}$ to $10^{-7}$. Performance was evaluated by measuring average accuracy across five FLOP measures (ranging from 2E+11 to 2E+15). For the baseline pattern task in Table 1 we choose a learning rate of 0.1 without performing hyperparameter optimisation as the experiment is simply to understand the behavior of LPN and not to optimize performance. For pattern OOD task we perform grid search for LPN and TTT over a dataset of from the strongly OOD task for 10 gradient steps, filtering out learning rates that do not decrease the test-time loss by at least 1.0%. We use these learning rates for the scaling specification size ablation also.
@@ Appendix G Architecture
In all our experiments, programs are defined in the input-output space of ARC-AGI, i.e., 2D grids whose cells can take 10 different values and have shapes $(n,m)$ with $n,m\in[1,30]$. We implement both the encoder and decoder as small transformers [Vaswani et al., 2017] specifically designed for this benchmark, in contrast to the more general large language models (LLMs) typically used [Wang et al., 2023].
@@
We model the input and output images as 2D grids, which we pad and flatten in a raster-scan fashion to form sequences of pixel values, each of size $30\times 30=900$ (see Figure 19). Each grid sequence is prefixed with shape information, namely two extra values for the number of rows and columns, resulting in sequences of 902 values. For both the Encoder and Decoder grid positions are encoded using RoPE [Su et al., 2024] [^1] https://github.com/crowsonkb/rope-flax for both row and column indices.
@@ G.1 Encoder
The encoder processes both the input and output grids from a given pair and returns a distribution of inferred program latents underlying the task. Specifically, it outputs the mean and diagonal covariance of a multivariate normal distribution from which latents can be sampled. Each grid sequence contains 902 values, and we add an extra CLS token for the output embedding, resulting in a total sequence length of 1805 for the encoder transformer. The encoder is implemented as a standard transformer [Vaswani et al., 2017] with pre-layer normalization [Baevski and Auli, 2018, Xiong et al., 2020] and multi-head attention. To incorporate identify the input from the output we add an embedding $\text{emb}(c)$, $c\in\{0,1\}$ is the channel index (0 for input, 1 for output). All 1800 color values (0 to 9), four shape values (1 to 30), and the CLS token are separately embedded into $\mathbb{R}^{H}$ using lookup tables. Padded tokens, determined by the shape values, are masked, and the sequence is fed into multiple transformer blocks, see Appendix E for hyperparameter details. In the encoder the attention mask is non-causal, allowing all non-padded tokens to attend to each other during encoding. The CLS embedding is passed through a layer normalization and two parallel dense layers to output the mean and diagonal log-covariance of the multivariate normal distribution over latents. Sampled program latents have dimension $d$, which may differ from the embedding dimension $H$.

### v3

@@ 6 Conclusion
Acknowledgments and Disclosure of Funding
We thank Google’s TPU Research Cloud (TRC) for supporting this research. We are also grateful to Nathan Grinsztajn, Natasha Butt, Levi Lelis, and Jessica Hu for their feedback on the early versions of the paper. Matthew Macfarlane is supported by the LIFT-project 019.011, which is partly financed by the Dutch Research Council (NWO). We also thank reviewers for their detailed feedback, which we have integrated into the final version of the paper.
@@ A.2 ARC-AGI
We evaluate LPN on the ARC-AGI benchmark by training it on the re-arc dataset [hodel2024addressingabstractionreasoningcorpus],a dataset designed to be in distribution relative to the ARC training set, previously used for training large language models and ARC specific transformers to solve ARC-AGI li2024tackling, akyurek2024surprising. Most previous works however expand beyond this dataset, usually synthetically adding additional programs [butt2024codeit], or crafted datasets. We restrict training to tasks from only this dataset with the goal of assessing generalization instead of closing a performance gap with more data. This also ensures LPN solely learns from core knowledge priors without data leakage from the evaluation set. The evaluation set is known to be significantly out of distribution and therefore represents a challenging generalization experiment. We train a 178M-parameter LPN with a 256-dim latent space for 100k steps (batch size 256) for 2 days on a TPU v4-32, see Appendix˜E for further details, in terms of data, this amounts to 51M I/O pairs. Leveraging insights from smaller scale experiments we make use of gradient in LPN during training. We train LPN in Grad 0 mode for 95k steps and then fine tune in Grad 1 for a further 5k steps. We outline our hyperparameter procedure for both LPN and TTT for ARC-AGI in Appendix˜E.
@@ B.1 LPN Training Extended
We provide an expanded version of Table˜1 with additional ablations for both the training and inference axes.
@@
Table 5: Ablation study of LPN training and inference methods on the Pattern task. Rows represent training methods: LPN Grad [N] for N gradient steps, LPN Grad [N] ⋆⋆ with gradient flow through latent optimization (analogous to meta-learning), and Sample [N] for N random samples. Columns represent inference methods: Grad [N] for N gradient ascent steps and Sample [N] for random search with N samples. Training was performed for 20k steps with 3 seeds, with performance reported as mean (standard deviation) over 3 runs. Bold values indicate the best performance per inference method.
@@ B.2 Adapting Out-Of-Distribution
We provide expanded results for the out-of-distribution (OOD) experiments outlined in Section˜5.5. We include results for varying levels of OOD starting with in-distribution, weak OOD, and strong OOD.
@@ B.3 Validating the Decoder
Training deep networks from scratch to solve ARC-like tasks has been challenging [li2024tacklingabstractionreasoningcorpus]. If it is the case that neural networks struggle even to learn to execute single programs this represents a significant bottleneck to training models from scratch on a broad distribution of programs. Therefore, before training LPN end-to-end, we conclusively show that our decoder architecture does not suffer from such a bottleneck, and can learn individual programs.
We show 5 of the 400 total tasks from the ARC-AGI training set, and for each of these tasks, we train a small LPN architecture of 800k parameters (except for the last task which required a bigger model with 8.7M parameters) on the corresponding task generator from re-arc [hodel2024addressingabstractionreasoningcorpus]. Specifically, we select the first five tasks from the arc-agi_training_challenges json file (007bbfb7, 00d62c1b, 017c7c7b, 025d127b, 045e512c) shown in figure Figure˜5.
@@
We evaluate both the distribution of re-arc generated tasks on which it is trained and on the true task from the ARC-AGI training set in figure Figure˜6. We show that for each task, the small LPN decoder-only model successfully learns individual programs and manages to solve the corresponding ARC-AGI task. Therefore, our model outperforms previously reported results in li2024tacklingabstractionreasoningcorpus, and concludes that our architecture does not suffer from a decoder bottleneck. Note that the encoder is not helpful in this experiment since the task is always the same. Our later results on ARC-AGI Section˜5.7 take this a step further and show that we can learn a single transformer architecture capable of executing over 270 programs in the ARC training dataset.
@@ B.4 Analyzing the Latent Space
To validate that the encoder is learning programs in its latent space, we design an even simpler task with small 4x4 grids that have 2x2 patterns. We train a small LPN model until convergence with a latent space of dimension 2 to easily visualize it in figure Figure˜7. Due to the simplicity of the task we train the model with mean training, i.e. no latent optimization. Because we are using a 2D Gaussian prior for the latent space, we can convert $\mathbb{R}^{2}$ to the unit square using the normal cumulative distribution function (CDF), and then plot on the unit square at coordinates $(x,y)$ the decoder’s output when conditioned by the latent $\textrm{CDF}(z)=(x,y)$, or equivalently, $z=(\textrm{PPF}(x),\textrm{PPF}(y))$ using the percent-point function (PPF). These results demonstrate the diversity and smoothness of the latent space, showing structure in terms of color patterns, which motivates performing gradient ascent latent optimization in more complex tasks. This shows that the latent space can encode a wide range of diversity in its latent space which is especially important for adapting to unseen patterns.
@@ B.6 Measuring Floating Point Operations
The LPN Sampling results can be understood as ablation of LPN without gradient-based search, but only sampling multiple latents from the encoder computing the decoder likelihood for each, avoiding the computational cost of backpropagating through the decoder. However, this efficiency gain comes at a significant cost in performance, as shown in the pattern task results Table˜1.
@@ B.9 ARC-AGI Solution Analysis
B.10 Investigating Composition
To evaluate the generalization capabilities of LPN to novel compositions of learned primitives, we construct simple test cases by combining two distinct operations from the training set that the model has only encountered separately during training. This setup probes whether LPN can flexibly integrate these primitives to produce the desired composite effect, a key indicator of its ability to handle systematic generalization beyond rote memorization.
Figure 17: Testing LPN on the composition of two operations previously only seen separately.
In Figure 17, we examine a specific instance where the first operation extracts the pattern within a bounding box, while the second replaces all black pixels with pink. The direct output of the encoder reveals overfitting to isolated training patterns: it correctly extracts and the pattern within the bounding box, but fails to convert black pixels to black. However, by performing gradient ascent in the latent space to refine the encoding, we can mitigate these errors. This optimization aligns the latent representation more closely with the composite objective, enabling the decoder to accurately render the full pink conversion across the extracted pattern. This demonstrates LPN’s potential for compositional reasoning, where latent-space adjustments unlock emergent behaviors not explicitly trained for.
Figure 18: Testing LPN on the composition of three operations.
In Figure 18, we examine the combination of three primitives: a gravity-type operation, a fill operation, and the use of a checkerboard pattern for the fill. While LPN can directly predict the gravity and fill components, it overfits to the fill operations seen during training. As a result, even with a gradient-based search, it is not able to resolve this error and apply the checkerboard pattern instead.
B.11 LPN ARC-AGI Inference Ablations
To dissect the contributions of key components to LPN’s performance on ARC-AGI, we performed ablations on various test-time inference strategies. These experiments isolate the roles of the program prior and gradient-based search, using the ARC-AGI evaluation set (out-of-distribution relative to Re-ARC). We fix the number of sampled latents during inference and vary the search budgets across $\{10,50,100,200,400\}$ steps, evaluating on a single seed.
We compare the following inference strategies:
  • Encoder Sampling. Samples multiple latents from the encoder distribution, evaluates the specification likelihood for each, and selects the highest-scoring one. This assesses structured gradient-based search against brute-force sampling near the encoder output.
  • Gaussian Init + Gradient Search. Employs the core LPN gradient-based search but initializes from a sample drawn from the Gaussian prior. This isolates the encoder’s contributions in providing strong initializations and enhancing overall performance.
  • Encoder Init + Local Search. Initializes from the encoder output but substitutes gradient updates with a mutation operator (continuous noise perturbations). This evaluates gradient-based optimization against unstructured local search, while permitting exploration beyond the encoder distribution.
  • Gaussian Init + Local Search. Pairs prior-based initialization with local search. This baseline quantifies the joint necessity of encoder-guided initialization and gradient direction.
  • Encoder Init + Gradient Search (LPN). The full LPN inference procedure.
Table 11 reports the performance (in %) across these strategies and budgets.
Table 11: Ablation results (%) on ARC-AGI evaluation for inference strategies under varying search budgets.
Inference Method 10 50 100 200 400
Encoder Sampling 9.12% 10.75% 10.13% 10.50% 10.62%
Gaussian Init + Gradient Search 1.38% 2.75% 5.37% 6.88% 10.75%
Encoder Init + Local Search 8.38% 9.50% 10.37% 10.75% 10.75%
Gaussian Init + Local Search 0.25% 0.25% 0.25% 0.75% 0.25%
Encoder Init + Gradient Search (LPN) 9.10% 13.25% 15.00% 15.50% 15.50%
These results reveal that repeated encoder sampling yields only marginal gains over a single sample (7.75%), underscoring the limitations of brute-force approaches. The gap between Gaussian Init + Gradient Search and Encoder Init + Gradient Search emphasizes the encoder’s dual role: seeding effective starting points for low budgets and amplifying gains at high budgets (e.g., +4.75% at 400 steps). Encoder Init + Local Search matches gradient search at low budgets (10 steps) but plateaus thereafter, highlighting gradient guidance’s superiority for deeper exploration. Gaussian Init + Local Search yields near-zero performance, confirming the indispensability of encoder initialization paired with directed optimization. In aggregate, Encoder Init + Gradient Search (LPN) dominates all baselines, validating the encoder-informed initialization and gradient-based refinement as pivotal to LPN’s efficacy on ARC-AGI.
@@ Appendix C LPN Algorithm
Below we outline two algorithms: first, LPN test-time inference (Algorithm˜1) and its mechanism for performing inductive inference. Second, we provide the full algorithm for LPN during training (Algorithm˜2).
@@
5:   Sample $z_{i}\sim q_{\phi}(z|x_{i},y_{i})$ $\triangleright$ Using the reparameterization trick
@@
10:     $z_{i}^{\prime}\leftarrow z_{i}^{\prime}+\alpha\cdot\nabla_{z}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{n}\log p_{\theta}(y_{j}|x_{j},z)|_{z=z_{i}^{\prime}}$ $\triangleright$ Optional stop-grad on the 2nd term
@@
14:  $\mathcal{L}\leftarrow\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}_{i}$ $\triangleright$ Total loss for all pairs
@@ Appendix D Variational Inference
Latent Program Networks (LPNs), when operated in grad 0 mode, perform a form of variational inference. Updating the latent representation using gradients from the encoder constitutes semi-amortized variational inference [kim2018semi, marino2018iterative].
In prior work, such as LEAPs trivedi2021learning, it is assumed during training that the full representation of a program $f$ is observable. In this setting, variational inference on programs can be conducted using a standard Variational Autoencoder (VAE) restricted to the program space. LEAPs introduce a form of variational inference via an Execution Evidence Lower Bound (ELBO), where function reconstruction is based on correctly executing the function for a given input rather than reconstructing the full program representation. The Execution ELBO is defined as:
@@ Appendix E Hyperparameters
For the ARC-AGI results, we performed a hyperparameter search over the learning rate for test time adaptation for both test-time tuning and LPN. Since these methods operate in different spaces, they are unlikely to require similar parameters, making it fairer to tune this parameter independently for each baseline. We validated on a held-out test set of unseen problems from the RE-ARC dataset. We searched over learning rates ranging from $10^{-1}$ to $10^{-7}$. Performance was evaluated by measuring average accuracy across five FLOP measures (ranging from 2E+11 to 2E+15). For the baseline pattern task in Table˜1 we choose a learning rate of 0.1 without performing hyperparameter optimisation as the experiment is simply to understand the behavior of LPN and not to optimize performance. For pattern OOD task we perform grid search for LPN and TTT over a dataset of from the strongly OOD task for 10 gradient steps, filtering out learning rates that do not decrease the test-time loss by at least 1.0%. We use these learning rates for the scaling specification size ablation also.
@@ Appendix G Architecture
In all our experiments, programs are defined in the input-output space of ARC-AGI, i.e., 2D grids whose cells can take 10 different values and have shapes $(n,m)$ with $n,m\in[1,30]$. We implement both the encoder and decoder as small transformers [vaswani2017attention] specifically designed for this benchmark, in contrast to the more general large language models (LLMs) typically used [wang2023codet5].
@@
We model the input and output images as 2D grids, which we pad and flatten in a raster-scan fashion to form sequences of pixel values, each of size $30\times 30=900$ (see Figure 21). Each grid sequence is prefixed with shape information, namely two extra values for the number of rows and columns, resulting in sequences of 902 values. For both the Encoder and Decoder grid positions are encoded using RoPE [su2024roformer]https://github.com/crowsonkb/rope-flax for both row and column indices.
@@ G.1 Encoder
The encoder processes both the input and output grids from a given pair and returns a distribution of inferred program latents underlying the task. Specifically, it outputs the mean and diagonal covariance of a multivariate normal distribution from which latents can be sampled. Each grid sequence contains 902 values, and we add an extra CLS token for the output embedding, resulting in a total sequence length of 1805 for the encoder transformer. The encoder is implemented as a standard transformer [vaswani2017attention] with pre-layer normalization [baevski2018adaptive, xiong2020layer] and multi-head attention. To incorporate identify the input from the output we add an embedding $\text{emb}(c)$, $c\in\{0,1\}$ is the channel index (0 for input, 1 for output). All 1800 color values (0 to 9), four shape values (1 to 30), and the CLS token are separately embedded into $\mathbb{R}^{H}$ using lookup tables. Padded tokens, determined by the shape values, are masked, and the sequence is fed into multiple transformer blocks, see Appendix˜E for hyperparameter details. In the encoder the attention mask is non-causal, allowing all non-padded tokens to attend to each other during encoding. The CLS embedding is passed through a layer normalization and two parallel dense layers to output the mean and diagonal log-covariance of the multivariate normal distribution over latents. Sampled program latents have dimension $d$, which may differ from the embedding dimension $H$.